# UniMoMo: Unified Generative Modeling of 3D Molecules for De Novo Binder Design

### v1

@@ Abstract
The design of target-specific molecules such as small molecules, peptides, and antibodies is vital for biological research and drug discovery. Existing generative methods are restricted to single-domain molecules, failing to address versatile therapeutic needs or utilize cross-domain transferability to enhance model performance. In this paper, we introduce Unified generative Modeling of 3D Molecules (UniMoMo), the first framework capable of designing binders of multiple molecular domains using a single model. In particular, UniMoMo unifies the representations of different molecules as graphs of blocks, where each block corresponds to either a standard amino acid or a molecular fragment. Based on these unified representations, UniMoMo utilizes a geometric latent diffusion model for 3D molecular generation, featuring an iterative full-atom autoencoder to compress blocks into latent space points, followed by an E(3)-equivariant diffusion process. Extensive benchmarks across peptides, antibodies, and small molecules demonstrate the superiority of our unified framework over existing domain-specific models, highlighting the benefits of multi-domain training.
@@ 3.1 Notations and Unified Representation
We represent a molecule as a graph of blocks ${\mathcal{G}}=({\mathcal{V}},{\mathcal{E}})$. ${\mathcal{V}}=\{({\bm{A}}_{i},\vec{{\bm{X}}}_{i})\mid{\bm{A}}_{i}\in{\mathbb{Z}}^{n_{i}},\vec{{\bm{X}}}_{i}\in\mathbb{R}^{n_{i}\times 3}\}$, where each block $i$ has $n_{i}$ atoms, with ${\bm{A}}_{i}$ and $\vec{{\bm{X}}}_{i}$ representing their element types and coordinates, respectively. =${\mathcal{E}}=\{({\bm{B}}_{i},{\bm{B}}_{ij})\mid i\neq j\}$, where ${\bm{B}}_{i}$ and ${\bm{B}}_{ij}$ denote intra-block and inter-block chemical bonds, respectively. For simplicity, we use matrix notation here, but the atoms within each block are unordered. To decompose molecules into blocks, we first identify all standard amino acids as individual blocks. In other cases, such as non-standard amino acids and small molecules, we leverage the principal subgraph algorithm (Kong et al., 2022) for decomposition, as illustrated in Figure 2. This algorithm begins with an atom-level graph and iteratively merges the frequent neighboring nodes into common building blocks such as benzene and indole (Appendix A). The resulting vocabulary of blocks, denoted by ${\mathbb{V}}$, comprises standard amino acids and the extracted principal subgraphs. Consequently, each block $i$ in the molecular graph can be assigned with a type $s_{i}\in{\mathbb{V}}$.
@@ 3.2 Iterative Full-Atom Variational AutoEncoder
where ${\bm{H}}_{i}^{t}$ and $\vec{{\bm{V}}}_{i}^{t}$ are the atomic hidden states and the motion vectors for each atom in block $i$. The time variable $t$ decreases from $1.0$ to $0.0$ with step size $\Delta t=0.1$ in this paper, totaling 10 iterations. At $t=1.0$, the initial coordinates $\vec{{\bm{X}}}_{i}^{1}$ are sampled from a Gaussian distribution ${\mathcal{N}}(\vec{{\bm{z}}}_{i},{\bm{I}})$ for each block $i$. The chemical bonds between atoms in different blocks are also predicted based on the spatial neighborhood (distance below 3.5Å) of atoms:
@@
where $b_{ij,pq}^{t}$ represents the type of chemical bond (e.g., single, double, triple, or none) between the $p$-th atom in block $i$ and the $q$-th atom in block $j$. Furthermore, ${\bm{h}}_{i,p}^{t}={\bm{H}}_{i}^{t}[p]$ and $\vec{{\bm{x}}}_{i,p}^{t}=\vec{{\bm{X}}}_{i}^{t}[p]$ denotes the hidden states and coordinates of the $p$-th atom in block $i$, respectively. $\operatorname{MLP}$ denotes a 3-layer multi-layer perceptron with SiLU activation (Hendrycks & Gimpel, 2016), the input of which is the summation of ${\bm{h}}_{i,p}^{t}$ and ${\bm{h}}_{j,q}^{t}$ to enable commutativity. Bond prediction is restricted to spatially neighboring atoms within 3.5Å, determined by the distance $d_{ij,pq}^{t}=\|\vec{{\bm{x}}}_{i,p}^{t}-\vec{{\bm{x}}}_{j,q}^{t}\|_{2}$. The reconstruction loss includes cross entropy (CE) for block types and bond types, as well as the mean square error (MSE) for motion vectors:
@@
where $p(\hat{s}_{i}),p(\hat{b}_{ij,pq}^{t})$, and $\hat{\vec{{\bm{V}}}}_{i,\text{gt}}^{t}$ denote the ground-truth block types, bond types, and motion vectors, respectively. To encourage better reconstruction of the interactions and local geometries, we also implement a pair-wise distance loss ${\mathcal{L}}_{\text{dist}}$ on the neighborhood for each atom. As the geometries are still in chaos on large time values, we only implement this loss when $t<0.25$:
@@
Both the encoder and the decoder are parameterized by an equivariant transformer (Jiao et al., 2024), with details in Appendix D. Training employs teacher forcing, where the structure module receives ground-truth atom types $\hat{{\bm{A}}}_{i}$ and intra-block bonds $\hat{{\bm{B}}}i$, and, with 50% probability, inter-block bonds $\hat{{\bm{B}}}{ij}$. During inference, full-atom structures and inter-block bonds are generated simultaneously and refined in an additional encoding-decoding cycle with the generated inter-block bonds also as inputs. Time values $t$ are randomly sampled at each training step, with the inputting coordinates $\vec{{\bm{X}}}_{i}^{t}$ interpolated between initial Gaussian-sampled coordinates $\vec{{\bm{X}}}_{i}^{1}$ and ground-truth coordinates $\vec{{\bm{X}}}_{i}^{0}$. Detailed training and sampling algorithms are provided in Appendix B.
@@ 3.3 Geometric Latent Diffusion Model
$\vec{{\bm{u}}}_{i}^{t}=\sqrt{\bar{\alpha}^{t}}\vec{{\bm{u}}}_{i}^{0}+(1-\bar{\alpha}^{t}){\bm{\epsilon}}_{i},$ (13)
@@
where $|{\mathcal{Z}}_{x}^{t}|$ indicates the number of nodes in the latent point cloud ${\mathcal{Z}}_{x}^{t}$. The overall training and sampling algorithms are included in Appendix C. It is worth noting that implementing a diffusion model in the latent space circumvents iterations over full-atom geometries, and thus tremendously reduce the computational cost for training and sampling.
@@ 4 Experiments
For ablation, we include two variants of UniMoMo in the tables throughout this section: one trained exclusively on domain-specific data (single) and the other on all datasets (all). Note that UniMoMo (all) employs the same model weights across benchmarks for different molecular binders, demonstrating its ability to generalize effectively across diverse molecular domains. Other ablations are in Appendix I. In the tables, the best one is marked in bold, with the second best underlined. For small molecules, the third best is also underlined following CBGBench (Lin et al., 2024d).
@@ 4.2 Antibody
Table 2: Results of recovery for antibody design.
@@
Table 3: Results of rationality for antibody design.
@@ 4.3 Small Molecule
Setup  We evaluate the results using the CBGBench framework (Lin et al., 2024d), which assesses models across four categories: substructure, chemical property, geometry, and interaction. Substructure metrics calculate JSD and MAE between the generated and reference distributions of atoms, rings, and functional groups. Chemical property metrics include QED, SA, LogP, and LPSK. Geometry metrics focus on JSD for bond lengths and angles, as well as atomic and molecular clash rates. Interaction metrics evaluate Vina energy in score, minimization, and dock modes, along with interaction type distributions between molecules and proteins. Models are ranked within each category, with ranks converted to scores using the formula $(N-\text{rank})$, where $N$ is the total number of models. The overall score is a weighted sum of scores across categories, the higher, the better. Detailed descriptions are available in Appendix F.
@@
Results  Table 4 presents the overall ranking scores, with detailed evaluations summarized in Tables 567, and 9 (Appendix G). UniMoMo achieves the best fidelity to substructure distributions (Table 5), excelling in JSD and MAE for atom types, rings, and functional groups, which is due to the block-level unified representation that effectively captures frequent motifs, such as aromatic rings and functional groups (Kong et al., 2022). The chemical property evaluation in Table 6 reveals the exceptional performance of our UniMoMo in generating drug-like (QED, Bickerton et al. (2012)) molecules with superior synthetic accessibility (SA, Ertl & Schuffenhauer (2009)). The geometry analysis in Table 7 highlights strength of UniMoMo in producing molecular binders with minimal steric clashes and well-distributed bond lengths and angles. In Table 9, UniMoMo exhibits competitive Vina docking scores and interaction profiles closely resembling the reference distribution, suggesting its capability of forming stable and meaningful interactions with target proteins. These results reflect higher structural rationalities, indicating that UniMoMo learns meaningful atom-level geometric constraints with iterative decoding in the full-atom autoencoder. As shown in Table 4, UniMoMo achieves the highest overall ranking score due to its consistent superiority across the comprehensive evaluation aspects. Importantly, the inclusion of multi-domain data significantly enhances performance in each aspect, indicating that UniMoMo is capable of learning a broader spectrum of property-structure relationships from all types of molecules. This could prevent biases seen in single-domain models, which helps the model generalize better and avoid overfitting to specific chemical spaces, therefore results in a model with better overall performance.
@@ 4.4 Different Binders for G Protein-Coupled Receptor
G protein-coupled receptors (GPCRs), the largest protein family in humans, are among the most prominent drug targets (Yang et al., 2021). We employ UniMoMo to generate peptide, antibody, and small molecule binders targeting the same binding site on a GPCR (PDB ID: 8U4R). Figure 3 shows UniMoMo produces candidates with strong in silico binding affinities, assessed by Rosetta (Alford et al., 2017) and Vina (Trott & Olson, 2010), and realistic geometries without requiring relaxation by physical forcefields. The showcased small molecule is particularly interesting, revealing two intriguing phenomena: 1) Interaction patterns from amino acids: The small molecule mimics the side chain of Arginine (orange box), one of the standard amino acids, to form hydrogen bonds with the target protein. 2) Local geometries from peptides and antibodies: To fit in the large pocket originally designed for antibodies, UniMoMo effectively generates wide-spanning scaffolds with amide connections (red boxes) typically seen in peptides and antibodies. The above findings highlight the strong generalization of our UniMoMo across diverse classes of molecular binders, with successful attempts to leverage cross-domain knowledge of molecular interactions and local geometries. Examples on other targets are included in Appendix H.
@@ Appendix B Algorithms for Training and Sampling with the Iterative Full-Atom Variational Autoencoder
Appendix C Algorithms for Training and Sampling with the Latent Diffusion Model

### v2

@@ Abstract
The design of target-specific molecules such as small molecules, peptides, and antibodies is vital for biological research and drug discovery. Existing generative methods are restricted to single-domain molecules, failing to address versatile therapeutic needs or utilize cross-domain transferability to enhance model performance. In this paper, we introduce Unified generative Modeling of 3D Molecules (UniMoMo), the first framework capable of designing binders of multiple molecular domains using a single model. In particular, UniMoMo unifies the representations of different molecules as graphs of blocks, where each block corresponds to either a standard amino acid or a molecular fragment. Subsequently, UniMoMo utilizes a geometric latent diffusion model for 3D molecular generation, featuring an iterative full-atom autoencoder to compress blocks into latent space points, followed by an E(3)-equivariant diffusion process. Extensive benchmarks across peptides, antibodies, and small molecules demonstrate the superiority of our unified framework over existing domain-specific models, highlighting the benefits of multi-domain training.
@@ 3.1 Notations and Unified Representation
We represent a molecule as a graph of blocks ${\mathcal{G}}=({\mathcal{V}},{\mathcal{E}})$. ${\mathcal{V}}=\{({\bm{A}}_{i},\vec{{\bm{X}}}_{i})\mid{\bm{A}}_{i}\in{\mathbb{Z}}^{n_{i}},\vec{{\bm{X}}}_{i}\in\mathbb{R}^{n_{i}\times 3}\}$, where each block $i$ has $n_{i}$ atoms, with ${\bm{A}}_{i}$ and $\vec{{\bm{X}}}_{i}$ representing their element types and coordinates, respectively. =${\mathcal{E}}=\{({\bm{B}}_{i},{\bm{B}}_{ij})\mid i\neq j\}$, where ${\bm{B}}_{i}$ and ${\bm{B}}_{ij}$ denote intra-block and inter-block chemical bonds, respectively. For simplicity, we use matrix notation here, but the atoms within each block are unordered. To decompose molecules into blocks, we first identify all standard amino acids as individual blocks. In other cases, such as non-standard amino acids and small molecules, we leverage the principal subgraph algorithm (Kong et al., 2022) for decomposition, as illustrated in Figure 2. This algorithm begins with an atom-level graph and iteratively merges the frequent neighboring nodes into common building blocks such as benzene and indole (Appendix A). The resulting vocabulary of blocks, denoted by ${\mathbb{V}}$, comprises standard amino acids and the extracted principal subgraphs. Consequently, each block $i$ in the molecular graph can be assigned with a type $s_{i}\in{\mathbb{V}}$. For better controllability, we introduce a binary prompt for each block, with 1 indicating the generation of natural amino acids (AAs), and 0 indicating either AAs or molecular fragments. During benchmarking for peptides and antibodies, all blocks are assigned a prompt of 1 to enforce AA-only generation. For small molecules, we set the prompt to 0, allowing flexible sampling from a broader set of molecular fragments.
@@ 3.2 Iterative Full-Atom Variational AutoEncoder
where ${\bm{H}}_{i}^{t}$ and $\vec{{\bm{V}}}_{i}^{t}$ are the atomic hidden states and the vector fields for each atom in block $i$. The time variable $t$ decreases from $1.0$ to $0.0$ with step size $\Delta t=0.1$ in this paper, totaling 10 iterations. At $t=1.0$, the initial coordinates $\vec{{\bm{X}}}_{i}^{1}$ are sampled from a Gaussian distribution ${\mathcal{N}}(\vec{{\bm{z}}}_{i},{\bm{I}})$ for each block $i$. The chemical bonds between atoms in different blocks are also predicted based on the spatial neighborhood (distance below 3.5Å) of atoms:
@@
where $b_{ij,pq}^{t}$ represents the type of chemical bond (e.g., single, double, triple, or none) between the $p$-th atom in block $i$ and the $q$-th atom in block $j$. Furthermore, ${\bm{h}}_{i,p}^{t}={\bm{H}}_{i}^{t}[p]$ and $\vec{{\bm{x}}}_{i,p}^{t}=\vec{{\bm{X}}}_{i}^{t}[p]$ denotes the hidden states and coordinates of the $p$-th atom in block $i$, respectively. $\operatorname{MLP}$ denotes a 3-layer multi-layer perceptron with SiLU activation (Hendrycks & Gimpel, 2016), the input of which is the summation of ${\bm{h}}_{i,p}^{t}$ and ${\bm{h}}_{j,q}^{t}$ to enable commutativity. Bond prediction is restricted to spatially neighboring atoms within 3.5Å, determined by the distance $d_{ij,pq}^{t}=\|\vec{{\bm{x}}}_{i,p}^{t}-\vec{{\bm{x}}}_{j,q}^{t}\|_{2}$. While potential inconsistency between predicted bonds and distances exists, they could be easily distinguished and discarded (Appendix G). The reconstruction loss includes cross entropy (CE) for block types and bond types, as well as the mean square error (MSE) for vector fields:
@@
where $p(\hat{s}_{i}),p(\hat{b}_{ij,pq}^{t})$, and $\hat{\vec{{\bm{V}}}}_{i,\text{gt}}^{t}$ denote the ground-truth block types, bond types, and vector fields, respectively. To encourage better reconstruction of the interactions and local geometries, we also implement a pair-wise distance loss ${\mathcal{L}}_{\text{dist}}$ on the neighborhood for each atom. As the geometries are still in chaos on large time values, we only implement this loss when $t<0.25$:
@@
Both the encoder and the decoder are parameterized by an equivariant transformer (Jiao et al., 2024), with details in Appendix F. Training employs teacher forcing, where the structure module receives ground-truth atom types $\hat{{\bm{A}}}_{i}$ and intra-block bonds $\hat{{\bm{B}}}i$, and, with 50% probability, inter-block bonds $\hat{{\bm{B}}}{ij}$. During inference, full-atom structures and inter-block bonds are generated simultaneously and refined in an additional encoding-decoding cycle with the generated inter-block bonds also as inputs. Time values $t$ are randomly sampled at each training step, with the inputting coordinates $\vec{{\bm{X}}}_{i}^{t}$ interpolated between initial Gaussian-sampled coordinates $\vec{{\bm{X}}}_{i}^{1}$ and ground-truth coordinates $\vec{{\bm{X}}}_{i}^{0}$. Detailed training and sampling algorithms are provided in Appendix B. We provide additional discussions and ablation studies on the design choices of the VAE in Appendix C.
@@ 3.3 Geometric Latent Diffusion Model
$\vec{{\bm{u}}}_{i}^{t}=\sqrt{\bar{\alpha}^{t}}\vec{{\bm{u}}}_{i}^{0}+\sqrt{1-\bar{\alpha}^{t}}{\bm{\epsilon}}_{i},$ (13)
@@
where $|{\mathcal{Z}}_{x}^{t}|$ indicates the number of nodes in the latent point cloud ${\mathcal{Z}}_{x}^{t}$. The overall training and sampling algorithms are included in Appendix D, with discussions on the equivariance in Appendix E. It is worth noting that implementing a diffusion model in the latent space circumvents iterations over full-atom geometries, and thus tremendously reduce the computational cost for training and sampling.
@@ 4 Experiments
For ablation, we include two variants of UniMoMo in the tables throughout this section: one trained exclusively on domain-specific data (single) and the other on all datasets (all). Note that UniMoMo (all) employs the same model weights across benchmarks for different molecular binders, demonstrating its ability to generalize effectively across diverse molecular domains. Other ablations are in Appendix H. In the tables, the best one is marked in bold, with the second best underlined. For small molecules, the third best is also underlined following CBGBench (Lin et al., 2024d).
@@ 4.2 Antibody
Table 2: Results of recovery for antibody design on CDR-H3.
@@
Table 3: Results of rationality for antibody design on CDR-H3.
@@ 4.3 Small Molecule
Setup  We evaluate the results using the CBGBench framework (Lin et al., 2024d), which assesses models across four categories: substructure, chemical property, geometry, and interaction. Substructure metrics calculate JSD and MAE between the generated and reference distributions of atoms, rings, and functional groups. Chemical property metrics include QED, SA, LogP, and LPSK. Geometry metrics focus on JSD for bond lengths and angles, as well as atomic and molecular clash rates. Interaction metrics evaluate Vina energy in score, minimization, and dock modes, along with interaction type distributions between molecules and proteins. Models are ranked within each category, with ranks converted to scores using the formula $(N-\text{rank})$, where $N$ is the total number of models. The overall score is a weighted sum of scores across categories, the higher, the better. Detailed descriptions are available in Appendix J.
@@
Results  Table 4 presents the overall ranking scores, with detailed evaluations summarized in Tables 567, and 8.UniMoMo achieves the best fidelity to substructure distributions (Table 5), excelling in JSD and MAE for atom types, rings, and functional groups, which is due to the block-level unified representation that effectively captures frequent motifs, such as aromatic rings and functional groups (Kong et al., 2022). The chemical property evaluation in Table 6 reveals the exceptional performance of our UniMoMo in generating drug-like (QED, Bickerton et al. (2012)) molecules with superior synthetic accessibility (SA, Ertl & Schuffenhauer (2009)). The geometry analysis in Table 7 highlights strength of UniMoMo in producing molecular binders with minimal steric clashes and well-distributed bond lengths and angles. In Table 8, UniMoMo exhibits competitive Vina docking scores and interaction profiles closely resembling the reference distribution, suggesting its capability of forming stable and meaningful interactions with target proteins. These results reflect higher structural rationalities, indicating that UniMoMo learns meaningful atom-level geometric constraints with iterative decoding in the full-atom autoencoder. As shown in Table 4, UniMoMo achieves the highest overall ranking score due to its consistent superiority across the comprehensive evaluation aspects. Importantly, the inclusion of multi-domain data significantly enhances performance in each aspect, indicating that UniMoMo is capable of learning a broader spectrum of property-structure relationships from all types of molecules. This could prevent biases seen in single-domain models, which helps the model generalize better and avoid overfitting to specific chemical spaces, therefore results in a model with better overall performance.
Model Vina Score Vina Min Vina Dock PLIP Interaction Rank
E IMP (%) E IMP (%) E IMP (%) MPBG (%) LBE JSDOA MAEOA JSDPP MAEPP
[…]
@@ 4.4 Different Binders for G Protein-Coupled Receptor
G protein-coupled receptors (GPCRs), the largest protein family in humans, are among the most prominent drug targets (Yang et al., 2021). We employ UniMoMo to generate peptide, antibody, and small molecule binders targeting the same binding site on a GPCR (PDB ID: 8U4R). Figure 3 shows UniMoMo produces candidates with strong in silico binding affinities, assessed by Rosetta (Alford et al., 2017) and Vina (Trott & Olson, 2010), and realistic geometries without requiring relaxation by physical forcefields. The showcased small molecule is particularly interesting, revealing two intriguing phenomena: 1) Interaction patterns from amino acids: The small molecule mimics the side chain of Arginine (orange box), one of the standard amino acids, to form hydrogen bonds with the target protein. 2) Local geometries from peptides and antibodies: To fit in the large pocket originally designed for antibodies, UniMoMo effectively generates wide-spanning scaffolds with amide connections (red boxes) typically seen in peptides and antibodies. The above findings highlight the strong generalization of our UniMoMo across diverse classes of molecular binders, with successful attempts to leverage cross-domain knowledge of molecular interactions and local geometries. Examples on other targets are included in Appendix K. Further analysis reveals an interesting self-adaptive mechanism in block size selection, which depends on the number of blocks designated for generation on the same binding site, with details in Appendix L.
@@ 5 Conclusion
Code Availability
The codes for data processing, model definition, training, testing, and evaluations are available at https://github.com/kxz18/UniMoMo.
@@ Appendix B Algorithms for Training and Sampling with the Iterative Full-Atom Variational Autoencoder
Appendix C Design Choices on the Iterative Full-Atom Variational AutoEncoder
C.1 Why Do We Need the Compressed Latent Space?
We incorporate a latent space primarily to facilitate the implementation of the diffusion model. In our unified representation, each graph node corresponds to a block, such as a residue or a chemical fragment, whose atom count varies depending on its type. As a result, the block coordinates $\vec{{\bm{X}}}_{i}\in\mathbb{R}^{n_{i}\times 3}$ and atom types ${\bm{A}}_{i}\in\mathbb{Z}^{n_{i}}$ are variable in size during denoising, making direct diffusion on these inputs infeasible due to the fixed-dimensional nature of standard diffusion models. To overcome this, we employ an all-atom variational autoencoder (VAE) that encodes each block into fixed-length latent vectors ${\bm{z}}_{i}$ and $\vec{{\bm{z}}}_{i}$, thereby enabling diffusion to operate efficiently and consistently in this fixed-dimensional continuous latent space.
C.2 Balancing KL Weights on ${\bm{z}}_{i}$ and $\vec{{\bm{z}}}_{i}$
Different KL weights (i.e., $\lambda_{1}$ and $\lambda_{2}$ in Eq.2) applied to the invariant (${\bm{z}}_{i}$) and the equivariant ($\vec{{\bm{z}}}_{i}$) latent variables shape the latent space in distinct ways. To evaluate whether the resulting space is suitable for generative modeling, we use the diffusion loss as a proxy, with a lower loss indicating a better latent space for diffusion. Larger KL weights promote smoother, more continuous latent spaces that benefit diffusion modeling, but excessive regularization may overly compress the representations, diminishing expressivity through information loss. Therefore, these weights must strike a balance to ensure the latent space is both smooth and sufficiently informative. As shown in the validation loss curves in Figure4, we choose the weight combination that yields the lowest validation loss.
Figure 4: Validation loss curves for the latent diffusion model trained on VAE latent spaces with varying KL weights, where $\lambda_{1}$ and $\lambda_{2}$ indicate the KL weights on the invariant and the equivariant latent variables, respectively, as in Eq. 2.
C.3 Ablation Studies on Extra Noises on $\vec{{\bm{z}}}_{i}$
We manually introduce additional noise to the latent coordinates $\vec{{\bm{z}}}_{i}$ before decoding to enhance robustness against errors introduced by the diffusion model. This strategy is particularly beneficial for small molecules, which often exhibit more complex inter-block connectivity than peptides or antibodies, where blocks are typically connected by peptide bonds. These more intricate structures are more sensitive to coordinate perturbations, making robustness especially important. While increasing the KL weight might seem like an alternative way to enforce smoother latent spaces, our current setting of 0.8 for $\vec{{\bm{z}}}_{i}$ already imposes strong regularization. Further increasing it would excessively compress the latent space and degrade the diffusion model’s performance, as discussed in Figure 4. As shown in Table 9, adding extra noise significantly improves the tolerance of the autoencoder to coordinate inaccuracies from diffusion, tremendously enhancing stability and affinity of the generated molecules.
Table 9: Ablation studies on extra noises and increased KL weights on latent coordinates $\vec{{\bm{z}}}_{i}$.
model vina(score only) vina(minimize) vina(dock)
w/o extra noise + $\lambda_{2}=0.8$4.62 -5.58 -7.24
w/o extra noise + $\lambda_{2}=1.0$4.42 -5.59 -7.09
w/ extra noise + $\lambda_{2}=0.8$5.72 -6.08 -7.25
Appendix D Algorithms for Training and Sampling with the Latent Diffusion Model